Part of $5000 is invested at 4% and the rest at 5%. The annual interest on the 5% was $124.00 more than the interest on the 4% investment. How much was invested at each rate?
Not from a text book, this was a handed out assignment.

Answer by scott8148(6628)

x(4%)+(5000-x)(5%)=x(4%)+(x(4%)+124)

.04x+250-.05x=.08x+124 ... 126=.09x ... 1400=x ... 3600=5000-x